OpenStudy (anonymous):

if three cards are dealtt from a deck of 52, what is the probabilty that the first card is a heart and the next to are diamonds? please help :)

OpenStudy (dumbcow):

13 hearts in a deck 13 diamonds in a deck 1st card: P(heart) = #hearts/#cards = 13/52 =1/4 2nd card: P(diamond) = #diamonds/#cards = 13/51 3rd card: P(diamond) = #diamonds/#cards = 12/50=6/25 The drawing of each card is an independent event so we multiply probabilities P(HDD) = (1/4)*(13/51)*(6/25)

OpenStudy (anonymous):

THANKK you (:

OpenStudy (dumbcow):

no problem

OpenStudy (anonymous):

(: